OpenStudy (anonymous):

Which shows this equation in standard form? 7 – 3(x – y) = 5x + 2 A. 3x – 8y = 5 B. 8x + 3y = 5 C. 8x – 3y = 5 D. 8x – 3y = –5

OpenStudy (anonymous):

@twinz1999 @imawesome101

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

standard form is 3y-8x=-5

OpenStudy (anonymous):

and im 100% sure

OpenStudy (anonymous):

So D is ur answer

OpenStudy (anonymous):

i dont see that one o d ok ty do u want a medal and a fan

OpenStudy (anonymous):

u dont see it because mine is not in the same order

OpenStudy (anonymous):

3y-8x=-5 is the same as 8x-3y=-5

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

relist the answer choices correctly

OpenStudy (anonymous):

i did they are copyed and pasted

OpenStudy (anonymous):

so there is only one -5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

TWINZ!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

heyyyyyyyyyyyyyy

OpenStudy (anonymous):

hows it goin

OpenStudy (anonymous):

i lost yesterday by a freckin lugnut

OpenStudy (anonymous):

Good, wbu? :)

OpenStudy (anonymous):

im good just been racin :)

OpenStudy (anonymous):

creppy face :D

OpenStudy (anonymous):

what unit and lesson r u in on math

OpenStudy (anonymous):

cool :)creepy face lol

OpenStudy (anonymous):

and im not sure